OpenStudy (anonymous):

The circumference of a circle is 15ft. What is the length of the radius? Use 3.14 for pi. Round your answer to the nearest tenth of a foot?

OpenStudy (anonymous):

The equation for circumference is 2*pi*r. You have pi and the circumference (= 15ft). So the equation becomes: 15 = 2*pi*r. Now solve for r.

OpenStudy (anonymous):

I'm confused :c

OpenStudy (anonymous):

Okay, we'll do this step-by-step. The equation for the circumference of a circle is given by C = 2*pi*r, where C is the circumference (given as 15ft), pi is, well, pi (given as 3.14 in the question), and r is the radius of the circle. Our goal is to find the radius. First we substitute everything we know into the equation for the circumference, so it becomes: 15 = 2*3.14*r Are you following so far? Now, we divide both sides of the equation by 2*3.14, so that the right-side of the equation becomes only 'r' - like so: 15/(3.14*2) = (2*3.14*r)/(2*3.14) 15/(3.14*2) = r Now simply divide 15 by the product of the (3.14*2) to get the required radius in feet. :)

OpenStudy (anonymous):

6.28/ 15 right?

OpenStudy (anonymous):

No, the other way! 15/6.28

OpenStudy (anonymous):

Ohh, Okay! Thank you!

OpenStudy (anonymous):

2.388535031847134?